OpenStudy (anonymous):

What is the remainder when 2^(495) is divided by 11? No calculator allowed. Please show technique

OpenStudy (freckles):

hmmm... maybe we can try to use fermat's little theorem

OpenStudy (freckles):

\[495=11(45) \\ 2^{495}=2^{11 \cdot 45 }\] then later we can use that 45=4(11)+1

OpenStudy (freckles):

recall fermat's little theorem is \[a^p \equiv a (\mod p)\] p is prime

OpenStudy (anonymous):

I literally have no idea what you are doing lol

OpenStudy (freckles):

have you ever heard of fermat's little theorem

OpenStudy (freckles):

\[2^{495} \equiv (2^{11})^{45}\equiv 2^{45} \text{ by fermat's little theorem } \\ 2^{45} \equiv2^{11(4)+1}\equiv(2^{11})^42^{1} \equiv (2)^4(2)\equiv =2^5 \equiv 32 \] do you what 32 mod 11 is?

OpenStudy (freckles):

by the way i used fermat's little theorem twice

OpenStudy (freckles):

\[2^{495} \equiv (2^{11})^{45}\equiv 2^{45} \text{ by fermat's little theorem } \\ 2^{45} \equiv2^{11(4)+1}\equiv(2^{11})^42^{1} \equiv (2)^4(2) \text{ by fermat's little theorem again }\\ (2)^4(2)=2^5 \equiv 32 \equiv ?\] the last step I'm asking you to do what 32 mod 11?

OpenStudy (freckles):

@liliy have you left me already?